OpenStudy (serenavdw):

between what two consecutive integers is negetive square root of 30

OpenStudy (anonymous):

Clearly \[\sqrt{25} < \sqrt{30} < \sqrt{36}\] Which can be simplified to \[5 < \sqrt{30} < 6\] Therefore it follows that \[-6 < -\sqrt{30} < -5\] And so -√30 lies between the consecutive numbers -6 and -5.

OpenStudy (serenavdw):

thanks everyone

OpenStudy (anonymous):

Your welcome